OpenStudy (anonymous):

howl to solve this exponential equation for a medal? 5 9x = 12.48

OpenStudy (anonymous):

take the log of both sides

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

this will give you 9xlog(5) = log(12.48)

OpenStudy (anonymous):

So what would be the next step?

OpenStudy (anonymous):

then would you divide them?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so now you have 9x=log(12.48)/log(5)

OpenStudy (anonymous):

what's the last step?

OpenStudy (swamijatin2):

http://www.tiger-algebra.com/drill/5$9x=12.48/

OpenStudy (anonymous):

Cancel the logs?

OpenStudy (anonymous):

No, you want to isolate the x. So you would divide both sides by 9

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

x=log(12.48)/(log(5)(9))

OpenStudy (anonymous):

Just plug that into your calculator

OpenStudy (anonymous):

I don't have a scientific calculator right now

OpenStudy (anonymous):

x=0.174

OpenStudy (anonymous):

oh ok thanks

OpenStudy (triciaal):